OpenStudy (anonymous):

Simplify the expression. √27x^5 + x√3 + 9√3x If the simplified expression is written in standard form, what is the leading coefficient?

OpenStudy (zzr0ck3r):

sqrt(27x^5) + xsqrt(3) + 9sqrt(3x)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Not sure exactly what is intended by "simplify," but these are possible ways to state the expression:\[\sqrt{27x^5} + x \sqrt{3} + 9\sqrt{3x}=3x^2 \sqrt{3x}+x \sqrt3 +9\sqrt{3x}=\sqrt3[(3x^2+9) \sqrt x+x]\]

OpenStudy (anonymous):

OOPS! Ran off the end... Last form has sqrt3 factored out from all terms, and sqrt x from two of them. Not really sure that qualifies as simplified, but it might be useful in some circumstances.

OpenStudy (anonymous):

THank you so much! i appreciate it

OpenStudy (anonymous):

Another way to express it might be\[\sqrt3 [(x^2+3)3 \sqrt x+x]\]